TRUE OR FALSE

1)A-Your investment rule is to only buy triple net lease properties that offer a positive NPV when the net operating income and sale of property is discounted back to the present at a 10% cost of capital. In other words, the discount rate you use in your calculations is 10%. Assume that you will pay the asking price. Your investment horizon is 15 years. At the end of 15 years you expect to sell the property for 30% more than you paid. Ignore taxes on capital gains.

The price is $5 million. NOI is $300,000 per year for 15 years.

You should invest in the following property?

B-Your investment rule is to only buy triple net lease properties that offer a positive NPV when the net operating income and sale of property is discounted back to the present at a 15% cost of capital. In other words, the discount rate you use in your calculations is 15%. Assume that you will pay the asking price. Your investment horizon is 15 years. At the end of 15 years you expect to sell the property for 30% more than you paid. Ignore taxes on capital gains.

The price is $5 million. NOI is $200,000 per year for 15 years.

You should invest in the following property?

C- Your investment rule is to only buy triple net lease properties that offer a positive NPV when the net operating income and sale of property is discounted back to the present at a 15% cost of capital. In other words, the discount rate you use in your calculations is 15%. Assume that you will pay the asking price. Your investment horizon is 15 years. At the end of 15 years you expect to sell the property for 30% more than you paid. Ignore taxes on capital gains.

The price is $2 million. NOI is $400,000 per year for 15 years.

You should invest in the following property?
Answer #1

Answer 1:

FALSE (You should NOT invest in the property)

Explanation:

Investment rule is to only buy triple net lease properties that offer a positive NPV when the net operating income and sale of property is discounted back to the present at a 10% cost of capital.

PV Factor for a One-Dollar Annuity Discounted at k Percent for n Periods = [1 - 1/(1 + k) n] / k

PV Factor for One Dollar Discounted at k Percent for n Periods = 1 / (1 + k) n

NPV = PV of annual NOI over 15 years + PV of sale price at the end year 15 - current price paid

= 300000 * (1 - 1/(1+ 10%) 15 ) / 10% + 5000000 * (1 + 30%) * 1/ (1 + 10%) 15 - 5000000

= -$1,162,127.83

As NPV is negative you should not invest in the property.

Answer 2:

FALSE (You should NOT invest in the property)

Explanation:

NPV = PV of annual NOI over 15 years + PV of sale price at the end year 15 - current price paid

= 200000 * (1 - 1/(1+ 15%) 15 )/ 15% + 5000000 * (1 + 30%) * 1/ (1 + 15%) 15 - 5000000

= -$3,031,711.83

As NPV is negative you should not invest in the property.

Answer 3:

TRUE (You should invest in the property)

Explanation:

NPV = PV of annual NOI over 15 years + PV of sale price at the end year 15 - current price paid

= 400000 * (1 - 1/(1+ 15%) 15 )/ 15% + 2000000 * (1 + 30%) * 1/ (1 + 15%) 15 - 2000000

= $658,473.70

As NPV is positive you should invest in the property.